# Tony Lindeberg

### v7

@@ 3 Results
To illustrate to what extent the distributions of the resultant will be influenced by receptive fields for different orders of spatial differentiation, the top row in Figure 10 shows histograms of the resultant accumulated for idealized models of simple cells of orders 1 and 2 as well as for the idealized model of complex cells, based on a quasi-quadrature combination of the responses from first-order and second-order simple cells. For generating these graphs, we have created a uniform distribution of the scale parameter ratio $\kappa$ over a logarithmic scale, over the interval $\kappa\in[1/\kappa_{\max},\kappa_{\max}]$ for the arbitrary choice of $\kappa_{\max}=8$. The bottom row in Figure 10 shows corresponding histograms of the resultant for third-order and fourth-order simple cells.
@@
As can be seen from these graphs, the histogram over the first-order simple cells is delimited by a maximum value around $R_{\max,1}\approx 0.7$, while the distributions for third-order and fourth-order simple cells are heavier for larger values of the resultant approaching $R\rightarrow 1$, and in line with the above arguments. The model used for computing histograms of the resultant for the idealized models of complex cells does, however, not very well reproduce the shape of the biologically obtained histogram, thus indicating that the model for the complex cells may be overly simplified [^3] In this context, it should be remarked that there is strong conceptual difference between the idealized models for simple cells vs. the idealized models of complex cells. The idealized model for simple cells has been determined in a theoretically principled manner from axiomatic derivations, and also been matched to biological measurements of simple cells, whereas the idealized model for complex cells has been chosen as an as straightforward way as possible for combining the responses of odd-shaped and even-shaped simple cells. for the purpose of reproducing the shape of the resultant histogram, see the discussion in Section 6 in (Lindeberg 2024a) for a number of suggestions concerning ways to extend that model.
@@
With ample reservation from the facts that: (i) in the current state, we have not principled biological arguments for choosing particular values of the parameters $\kappa_{\min}$ and $\kappa_{\max}$ that determine the range of the scale parameter ratio $\kappa$, where different choices of these parameters may affect the shapes of the combined histograms of the resultant $R$, (ii) the choice of a logarithmic distribution over the scale parameter $\kappa$ does, as previously mentioned, neglect any co-dependency with respect to the distribution over the orientation angle $\varphi$, (iv) the assumption about equal numbers of receptive fields for the different orders of spatial differentiation may not necessarily hold in reality, and additionally expressing reservation from the fact that (v) our computations of the resultant $R$ for the receptive fields are based on idealized noise free model, while there additionally could be sources to noise in the biological experiments as well as modelling errors between the receptive fields of the actual biological neurons and our idealized receptive fields, as can be seen from these results, these combined histograms give rise to a bump in the histograms for lower values of the resultant $R$, in qualitative agreement with the biological results by Goris et al. (2015) and as reproduced in Figure 7. Furthermore, to obtain something that would look like a small bump for larger values of the resultant $R$, the modelling situation with receptive fields up to order 4 would give a closer similarity to the biological results by Goris et al. (2015) compared to the model based on receptive fields up to order 2.
@@
From this analysis, it thus appears as if (i) the histogram of the resultant of the simple cells in the biological experiments et al. (2015) could be rather well explained by the receptive fields in the primary visual cortex of Macaque monkeys having a variability over the degree of elongation, and that additionally (ii) the non-uniform nature of the experimentally obtained histogram of the resultant $R$ for the simple cells could be explained better by assuming that receptive fields should be present up to a spatial differentiation order up to 4 than up to a spatial differentiation order of 2.
@@
3.1.4 Summary of the interpretations of the biological experiments by Nauhaus et al. (2008) and by Goris et al. (2015)
@@
With regard to the working hypothesis that we set out to investigate the possible validity of, we can conclude that the experimental results by both Nauhaus et al. (2008) and by Goris et al. (2015) are, in combination with the theoretical results in Sections 2.2 and A.1, concerning a direct connection between the degree of elongation of a receptive field with the orientation selectivity of the receptive field, clearly consistent with an expansion over the degree of elongation of the receptive field shapes in the primary visual cortex.
@@
Figure 12: Orientation map in the primary visual cortex of cat, as recorded by Koch et al. (2016) (OpenAccess), with the orientation preference of the receptive fields encoded in terms of colours, and demonstrating that the visual cortex performs an explicit expansion of the receptive field shapes over spatial image orientations. A working hypothesis in the paper concerns to investigate whether the primary visual cortex could additionally perform an expansion over the eccentricity or the elongation of the spatial components of the receptive fields. One possible way of performing such an additional expansion, is over the spatial covariance matrices $\Sigma$ of the affine Gaussian derivative kernels according to (5), and as illustrated in Figure 13, although that illustration would additionally need to be complemented by an identification of opposite image orientations, as, for example, can be achieved by a mapping to the double angle $\varphi\mapsto 2\varphi$, as well as a possible adjustment of the second degree of freedom, in how the variability of the two scale parameters in the affine Gaussian derivative model, beyond a variability over their ratio, varies from the center to the periphery in that illustration.
First-order affine Gaussian derivative kernels
@@ 3.2 Explicit testable hypotheses for biological experiments
If additionally, reconstructions of the receptive field shapes could be performed for the receptive fields probed during such a systematic investigation of the difference in response characteristics with the distance from the pinwheel centers, and if the receptive fields could additionally be reasonably well modelled according to the generalized Gaussian model for receptive fields studied and used in this paper, it would be interesting to investigate if the shapes of the affine Gaussian components of these receptive fields would span a larger part of the affine group, than the span over mere image orientations, as already established in the orientation maps of the visual cortex, as characterized by Bonhoeffer and Grinvald (1991), Blasdel (1992) and others, see Figure 13 for an illustration.
@@ 3.3 Quantitative measurements for detailed characterization
If the above theoretically motivated biological hypotheses could be investigated experimentally, and if the above quantitative measurements of receptive field characteristics could be performed, it could be judged if the prediction from the presented theoretical analysis about a systematic variability in receptive field eccentricity, with a possible relationship to the pinwheel structure, could be either experimentally supported or rejected. In a corresponding manner, such a judgement could also answer if the receptive fields in the primary visual cortex could be regarded as spanning a larger part of the affine group, than an expansion over mere rotations in the image domain.
4 Discussion
@@
By comparing the theoretical analysis of the orientation selectivity properties of the affine Gaussian derivative model (Section 2.2) with the experimental results from Nauhaus et al. (2008) on broadly vs. sharply tuned visual neurons (Figure 6) and from Goris et al. (2015) on rather uniform distributions of the resultant values from orientation selectivity curves (Figure 7), we find potential support for one of the dimensions of variability in a biological hypothesis formulated in Lindeberg (2023), stating that the family of receptive field shapes should span the degrees of freedom in the natural geometric image transformations. This potential support rests on the assumption, that it should be unlikely for the population of receptive fields to show strong variability in orientation selectivity, without also showing similar variability in eccentricity or elongation.
Such an assumption-based reasoning would then specifically imply indirect support for the hypothesis that the receptive field shapes should span a sufficiently wide range of ratios between the scale parameters in the directions perpendicular to versus parallel with the preferred orientation of the receptive field, to support affine covariance of the family of visual receptive fields.
@@
Without explicitly relying on expressing such an explicit assumption, regarding whether the visual receptive fields in the primary visual cortex could be well modelled by affine Gaussian derivative based receptive fields, we can, however, firmly state that the biological measurements performed by Nauhaus et al. (2008) and by Goris et al. (2015) are, in combination with the theoretical results summarized in Section 2.2, consistent with the hypothesis that the receptive fields should span a variability in the eccentricity of the receptive fields. In this respect, the measurements that demonstrate a strong variability in orientation selectivity would specifically be consistent with the theoretically based biological hypothesis formulated in (Lindeberg 2023), that the receptive fields in the primary visual cortex should span the variability of receptive field shapes under spatial affine transformations.
@@
Furthermore, from the results presented in Section 3.1.3, that the orientation selectivity histograms accumulated by Goris et al. (2015 appear to be better explained by a population of receptive fields up to order 4 than by receptive fields up to order 2, one may speculate if it would be appropriate to put additional focus of neurophysiological experiments on exploring the presence of visual receptive fields of higher order, compared to the receptive fields of lower order, that are more commonly reported in the neurophysiological literature.
@@
If we would apply a similar type of assumption-based logical reasoning to the pinwheel structure in the primary visual cortex, then such a reasoning, based on the results by Nauhaus et al. (2008), that the orientation selectivity appears to vary strongly from the centers of the pinwheels towards the periphery, would imply that the pinwheel structure in the visual cortex would, beyond an explicit expansion over image orientations, also comprise an explicit expansion over the eccentricity or the degree of elongation of the receptive fields. Based on these predictions, we propose to consider explicit dependencies on a variability in the eccentricity of the receptive fields, when modelling the pinwheel structures in the primary visual cortex.
@@
Strictly, and formally, the results from such logical inference could, however, only be regarded as theoretical predictions, to generate explicit hypothesis concerning the distribution of receptive field characteristics in these respects. To raise the question of determining if these theoretical predictions would firmly hold in reality, we propose that the testable explicit biological hypotheses formulated in Section 3.2 could be used to, in neurophysiological experiments, either verify or reject the overall hypothesis, concerning possible variabilities in the eccentricity of the receptive fields in the primary visual cortex of higher mammals, as well as hypotheses about possible connections between such variabilities in the eccentricity or the elongation and other receptive field characteristics, in particular in relation to the pinwheel structure in the primary visual cortex of higher mammals. Furthermore, if those hypothesis would hold, then the proposed quantitative measurements formulated in Section 3.3 could be used, to characterise how a possible variability in the eccentricity of the receptive field could be related to other receptive field characteristics, including the pinwheel structure in the visual cortex.
@@
Concerning possible limitations in the hypothetical reasoning stages used for possible logical inference and for formulating the explicit biological hypotheses above, based on explicitly stated assumptions regarding whether the biological receptive fields could be reasonably well modelled by affine Gaussian derivative based receptive fields, to be able to draw possible further conclusions, the possible validity of those hypothetical logical reasoning stages could, however, break down, if there would be other external factors, not covered by the theoretical model, that could also strongly influence the orientation selectivity of the receptive fields. The possible applicability of the hypothetical logical reasoning stages above thus, ultimately, depends on the possible agreement between the model and biological data, and can only be taken further by performing more detailed model fitting (not performed here, because of lack of access to the data by Nauhaus et al. (2008) as well as lack of access to data with receptive field recordings over a sufficiently large population of visual neurons in the primary visual cortex) and/or performing complementary neurophysiological experiments, to ultimately judge if the theoretically based predictions, stated more explicitly in Section 3.2, would be applicable to actual biological neurons.
@@
Furthermore, while the treatment in this paper has been concerned with species that have orientation maps and a pinwheel structure, a very interesting follow-up question would then also concern if corresponding results would extend to species that do not have structured orientation maps or a pinwheel structure, and then specifically which such species. For analyzing that topic in detail, further results from biological experiments would, however, be necessary, why we leave that topic to future work.
4.1 Relations to a corresponding analysis based on Gabor models
@@
In Lindeberg (2024a) Section 5, it is shown that if we couple the spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$ in that model according to $\sigma_{2}=\kappa\,\sigma_{1}$ for $\kappa>1$, then larger values of the scale parameter $\kappa$ will lead to more narrow orientation selectivity properties. There is, however, also another degree of freedom in the affine Gabor model, implying that variations in the angular frequency $\nu$ may also strongly affect the orientation selectivity properties of the receptive fields.
Specifically, comparing the explicit expressions for the orientation selectivity curves derived for the affine Gabor model in Lindeberg (2024a)
@@
Additionally, we have in Lindeberg (2024a) shown that the orientation selectivity curves for the joint spatio-temporal receptive field models (2.2.1) and (11) up to order 2 lead to similar orientation selectivity curves (16)–(18) as for the corresponding purely spatial models (5) and (8). In this respect, the analysis based on the generalized Gaussian derivative model for visual receptive fields has a much wider domain of validity, compared to a corresponding analysis based on the affine Gabor model for visual receptive fields, since its results hold also for the more realistic domain of joint spatio-temporal receptive fields.
4.2 Relations to more detailed and fine-grained models of the primary visual cortex

### v10

@@ 3.3.3 Orientation selectivity curves for the idealized receptive field models
To illustrate to what extent the distributions of the resultant will be influenced by receptive fields for different orders of spatial differentiation, Figure 9 shows histograms of the resultant accumulated for idealized models of simple cells of orders 1, 2, 3 and 4 as well as for the idealized model of complex cells. For generating these graphs, we have created a uniform distribution of the scale parameter ratio $\kappa$ over a logarithmic scale, over the interval $\kappa\in[1/\kappa_{\max},\kappa_{\max}]$ for the arbitrary choice of $\kappa_{\max}=8$.
@@
As can be seen from these graphs, the histogram over the first-order simple cells is delimited by a maximum value around $R_{\max,1}\approx 0.7$, while the distributions for third-order and fourth-order simple cells are heavier for larger values of the resultant approaching $R\rightarrow 1$. The model used for computing histograms of the resultant for the idealized models of complex cells does, however, not very well reproduce the shape of the biologically obtained histogram, thus indicating that the model for the complex cells may be overly simplified [^8] In this context, it should be remarked that there is strong conceptual difference between the idealized models for simple cells vs. the idealized models of complex cells. The idealized model for simple cells has been determined in a theoretically principled manner from axiomatic derivations, and also been matched to biological measurements of simple cells, whereas the idealized model for complex cells has been chosen as an as straightforward way as possible for combining the responses of odd-shaped and even-shaped simple cells. for the purpose of reproducing the shape of the resultant histogram. See the discussion in Section 6 in (Lindeberg 2025) for a number of suggested ways to extend that model.
@@
With ample reservation from the fact that this theoretical analysis is conceptually simplified in a number of ways, [^9] The reservation that we have to state concerning this theoretical modelling of the resultant histograms accumulated by by Goris et al. (2015), however, is that the analysis is simplified in the following 5 major ways: (i) In the current state, we do not have any principled biological arguments for choosing particular values of the parameters $\kappa_{\min}$ and $\kappa_{\max}$ that determine the range of the scale parameter ratio $\kappa$, where different choices of these parameters may affect the shapes of the combined histograms of the resultant $R$. (ii) The choice of a logarithmic distribution over the scale parameter $\kappa$ does, as previously mentioned, neglect any co-dependency with respect to the distribution over the orientation angle $\varphi$. (iii) The experimental conditions used when accumulating the orientation selectivity curves for biological neurons may not be identical to the conditions under which the theoretically derived orientation selectivity curves have been derived. (iv) The assumption about equal numbers of receptive fields for the different orders of spatial differentiation may not necessarily hold in reality. (v) Our computations of the resultant $R$ for the receptive fields are based on an idealized noise free model, while there additionally could be sources of noise in the biological experiments as well as modelling errors between the receptive fields of the actual biological neurons and our idealized receptive fields. as can be seen from these results, the combined histograms give rise to a bump in the histograms for lower values of the resultant $R$, in qualitative agreement with the biological results by Goris et al. (2015) and as reproduced in Figure 6. Furthermore, to obtain something that would look like a small bump for larger values of the resultant $R$, the modelling situation with receptive fields up to order 4 gives a closer similarity to the biological results by Goris et al. (2015) compared to the model based on receptive fields up to order 2.
Assuming that the modelling simplifications do not significantly affect the qualitative nature of the results, from the presented analysis it thus appears as if:
@@
  • The histograms of the resultant of the simple cells in the biological experiments et al. (2015) could be rather well explained by the receptive fields in the primary visual cortex of Macaque monkeys having a variability over the degree of elongation.
  • The non-uniform nature of the experimentally obtained histogram of the resultant $R$ for the simple cells could be explained better by assuming that receptive fields should be present up to a spatial differentiation order up to 4 than up to a spatial differentiation order of 2.
@@
4.1.4 Summary of the interpretations and modelling results of the biological experiments by Nauhaus et al. (2008) and by Goris et al. (2015)
@@
To conclude, the biological results by Nauhaus et al. (2008), Goris et al. (2015) and Wilson et al. (2016) are clearly consistent with an expansion over the degree of elongation of the receptive field shapes in the primary visual cortex.
@@ 4.2 Explicit testable hypotheses for biological experiments
Figure 11: Orientation map in the primary visual cortex of cat, as recorded by Koch et al. (2016) (OpenAccess), with the orientation preference of the receptive fields encoded in terms of colours, and demonstrating that the visual cortex performs an explicit expansion of the receptive field shapes over spatial image orientations.
First-order affine Gaussian derivative kernels
If additionally, reconstructions of the receptive field shapes could be performed for the receptive fields probed during such a systematic investigation of the difference in response characteristics with the distance from the pinwheel centers, and if the receptive fields could additionally be reasonably well modelled according to the generalized Gaussian model for receptive fields studied and used in this paper. Then, it would be interesting to investigate if the shapes of the affine Gaussian components of these receptive fields would span a larger part of the affine group, than the span over mere image orientations, as already established in the orientation maps of the visual cortex, as characterized by Bonhoeffer and Grinvald (1991), Blasdel (1992) and others, see Figure 11 for an illustration. A working hypothesis in the paper concerns to investigate whether the primary visual cortex could additionally perform an expansion over the eccentricity or the elongation of the spatial components of the receptive fields.
@@ 4.3 Quantitative measurements for detailed characterization
If the above theoretically motivated biological hypotheses could be investigated experimentally, and if the above quantitative measurements of receptive field characteristics could be performed. Then, it could be judged if the prediction from the presented theoretical analysis about a systematic variability in receptive field eccentricity, with a possible relationship to the pinwheel structure, could be either experimentally supported or rejected. In a corresponding manner, such a judgement could also answer if the receptive fields in the primary visual cortex could be regarded as spanning a larger part of the affine group, than an expansion over mere rotations in the image domain.
5 Relations to other types of models of visual receptive fields
Concerning the analysis presented in the previous section, it should be emphasized that it constitutes a theoretical modelling step at a functional mathematical level. The complementary theoretical explanation in terms of a partial affine covariance property regarding the degree of freedom corresponding to non-uniform scaling transformation does furthermore operate at a coarse level of abstraction, regarding theoretically desirable properties of an idealized vision system.
Hence, this theoretical explanation should not be regarded as in conflict with other possible theoretical explanations in terms of learning of spatial receptive fields by sparse coding, which may also lead to learned receptive fields with different degrees of elongation (Olshausen and Field 1996, 1997, Rehn and Martin 2007, Zylberberg et al. 2011, King et al. 2013). Similarly, our conceptual explanation is neither in conflict with a more fine-grained computational explanation of varying receptive field shapes in terms of different “On” or “Off” responses within the receptive fields, as proposed by Martinez et al. (2005) and neurophysiologically investigated by Kremkow et al. (2016) and Jansen et al. (2019).
Instead, our proposed explanation is that the notion of affine covariance or more generally covariance to geometric image transformations may be an essential factor in the development of biological vision systems. As natural images are formed in the retinas of visual observers, the statistical properties of the image data that reach the visual sensor will be strongly influenced by the structure of the natural image transformations, when 3-D objects are projected to 2-D image data from different viewing positions, viewing directions and relative motions.
The result of learning receptive fields in a biological vision agent, should thereby be strongly influenced by the structure of the natural image transformations. For a biological creature, who depends strongly on the visual perception for its survival, it does therefore seem plausible that there could be an evolutionary pressure for the biological organism to develop its vision system to be well adapted to the influence of the natural image transformations that arise when observing a dynamic world. Beyond the treatment of whether biological vision has actually developed a variability in the eccentricity of the receptive fields, our theory offers a theoretical explanation of this property, solidly grounded in a principled formal theory of visual receptive fields.
Thus, one could expect that when specific learning approaches, such as the sparse coding mechanisms considered by Olshausen and Field (1996, 1997), Rehn and Martin (2007), Zylberberg et al. (2011) and King et al. (2013), are exposed to natural image data that contain the variabilities in image structures generated by geometric image transformations, it would be natural for such learning approaches to lead to learned receptive fields with different eccentricities. Specifically, one could then also conceive that different spatial shapes of such receptive fields to be biologically implemented in terms of different contributions form “On” and “Off” subregions, as considered by Martinez et al. (2005), Kremkow et al. (2016) and Jansen et al. (2019).
In this respect, the presented theory in this paper thus offers a theoretical explanation of these phenomena at a more general and functional level of abstraction, by assuming an idealized vision system that adapts its processing to the inherent structures of natural image structures as influenced by natural image transformations. The property that the receptive fields in the primary visual cortex should have a variability in their eccentricity is then a direct consequence of this general assumption, via the assumption of affine covariance. This could then also conceptually explain other possible explanations in terms of more explicit learning mechanisms or specific neurophysiological mechanisms.
In Appendices A and B we additionally: (i) give relations to a corresponding analysis based on Gabor models of the visual receptive fields, and (ii) describe relations to more detailed and fine-grained models of the primary visual cortex.
6 Summary and discussion
@@
We have compared results from theoretical analysis of the orientation selectivity properties of the affine Gaussian derivative model (Section 3.3) with experimental results by Nauhaus et al. (2008) on broadly vs. sharply tuned visual neurons (Figure 5) and by Goris et al. (2015) on rather uniform distributions of the resultant values from orientation selectivity curves (Figure 6). Thereby, we have found potential support for one of the dimensions of variability in a biological hypothesis formulated in Lindeberg (2023), stating that the family of receptive field shapes ought to span the degrees of freedom in the natural geometric image transformations. This potential support rests on the assumption, that it should be unlikely for the population of receptive fields to show strong variability in orientation selectivity, without also showing similar variability in eccentricity or elongation.
@@
Without explicitly relying on expressing such an explicit assumption, regarding whether the visual receptive fields in the primary visual cortex could be well modelled by affine Gaussian derivative based receptive fields, we can, however, firmly state that the biological measurements performed by Nauhaus et al. (2008) and by Goris et al. (2015) are, in combination with the theoretical results summarized in Section 3.3, consistent with the hypothesis that the receptive fields should span a variability in the eccentricity of the receptive fields, and more widely consistent with the hypothesis about affine covariant receptive fields.
@@
Furthermore, from the results presented in Section 4.1.3, that the orientation selectivity histograms accumulated by Goris et al. (2015) appear to be better explained by a population of receptive fields up to order 4 than by receptive fields up to order 2, one may speculate if it would be appropriate to put additional focus of neurophysiological experiments on exploring the presence of visual receptive fields of higher order, compared to the receptive fields of lower order more commonly reported in the neurophysiological literature.
@@
If we apply a similar type of assumption-based logical reasoning to the pinwheel structure in the primary visual cortex, then such a reasoning, based on the results by Nauhaus et al. (2008) and by Wilson et al. (2016), that the orientation selectivity appears to vary strongly from the centers of the pinwheels towards the periphery. Then, this implies that the pinwheel structure in the visual cortex would, beyond an explicit expansion over image orientations, also comprise an explicit expansion over the the degree of elongation of the receptive fields. Based on these predictions, we propose to consider explicit dependencies on a variability in the eccentricity of the receptive fields, when modelling the pinwheel structure in the primary visual cortex.
@@
Strictly, and formally, the results from such logical inference could, however, only be regarded as theoretical predictions, to generate explicit hypothesis concerning the distribution of receptive field characteristics in these respects. To raise the question of determining if these theoretical predictions would firmly hold in reality, we propose that the testable explicit biological hypotheses formulated in Section 4.2 could be used to, in neurophysiological experiments, either verify or reject the overall hypothesis, concerning possible variabilities in the eccentricity of the receptive fields in the primary visual cortex of higher mammals. These predictions could also be used to explore hypotheses about possible connections between such variabilities in the eccentricity or the elongation and other receptive field characteristics, in particular in relation to the pinwheel structure in the primary visual cortex of higher mammals.
Notably, the overall analysis in the paper is carried based on a framework that reflects functional properties of an idealized vision system, in terms of covariance under geometric image transformations, and using mathematical analysis as the main tool. Thereby, the analysis in the paper is performed without any extensive numerical simulations of explicit neuron models, which would otherwise have to contain many parameters, some of which commonly are unobservable and would have to be estimated indirectly from data.
@@
Concerning possible limitations in the hypothetical reasoning stages used for possible logical inference and for formulating the explicit biological hypotheses above, the logical reasoning based on connections between the eccentricity and the degree of elongation of the receptive fields depend on explicitly stated assumptions regarding whether the biological receptive fields could be reasonably well modelled by affine Gaussian derivative based receptive fields. To be able to draw possible further conclusions, the possible validity of those hypothetical logical reasoning stages could, however, break down, if there would be other external factors, not covered by the theoretical model, that could also strongly influence the orientation selectivity of the receptive fields. The possible applicability of the hypothetical logical reasoning stages above thus, ultimately, depends on the possible agreement between the model and biological data, and can only be taken further by performing complementary neurophysiological experiments, to ultimately judge if the theoretically based predictions, stated in Section 4.2, would be applicable to actual biological neurons.
@@
Furthermore, while the treatment in this paper has been concerned with species that have orientation maps and a pinwheel structure, a very interesting follow-up question would then also concern if corresponding results would extend to species that do not have structured orientation maps or a pinwheel structure, and then specifically which such species. For example, Niell and Stryker (2008) have shown that the mouse visual cortex has developed mechanisms of orientation selectivity, although not having pinwheel structures. For analyzing that topic in detail, further results from biological experiments and on other species would, however, be necessary, why we leave that topic to future work.
Appendix
Appendix A Relations to a corresponding analysis based on Gabor models
@@
In Lindeberg (2025) Section 5, it is shown that if we couple the spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$ in that model according to $\sigma_{2}=\kappa\,\sigma_{1}$ for $\kappa>1$, then larger values of the scale parameter $\kappa$ will lead to more narrow orientation selectivity properties. There is, however, also another degree of freedom in the affine Gabor model, implying that variations in the angular frequency $\nu$ may also strongly affect the orientation selectivity properties of the receptive fields.
Specifically, comparing the explicit expressions for the orientation selectivity curves derived for the affine Gabor model in Lindeberg (2025)
@@
Additionally, we have in Lindeberg (2025) shown that the orientation selectivity curves for the joint spatio-temporal receptive field models (3.3.1) and (11) up to order 2 lead to similar orientation selectivity curves (16)–(18) as for the corresponding purely spatial models (5) and (8). In this respect, the analysis based on the generalized Gaussian derivative model for visual receptive fields has a much wider domain of validity, compared to a corresponding analysis based on the affine Gabor model for visual receptive fields, since its results hold also for the more realistic domain of joint spatio-temporal receptive fields.
Appendix B Relations to more detailed and fine-grained models of the primary visual cortex